Opened 4 years ago

Closed 3 years ago

#590 closed flight processing (fixed)

RG13/08, flight day 234a/2015, Cambridge

Reported by: lah Owned by:
Priority: immediate Milestone: 2015 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by lah)

Data location: ~arsf/arsf_data/2015/flight_data/uk/RG13_08-2015_234a_Cambridge

Data arrived from ARSF via USB on 27/08/2015.

Scientific objective: Measuring ash die back.

Priority: Unknown

PI: David Coomes

Flown on same day as Monks Wood (RG13/08) and Alconbury(GB15/00).

Sensors:

  • Fenix
  • Leica LIDAR
  • FW Leica LIDAR
  • RCD

Change History (34)

comment:1 Changed 4 years ago by lah

  • Description modified (diff)

comment:2 Changed 4 years ago by lah

Unpacking

All extra files have been split out except from RCD, which is duplicated for 234a, 234b and 234c. Please only process RCD for one of these projects, then split once processed.

comment:3 Changed 4 years ago by dac

Fenix data needs processing for this (at least up to finding SCT values) as a priority to check everything is working OK with the instrument - in particular the navigation data.

comment:4 Changed 4 years ago by gej

Navigation Processing Started

Lat: 52 11 07.33060
Long: -0 06 44.97300
ell.Height 119.956

Last edited 4 years ago by gej (previous) (diff)

comment:5 Changed 4 years ago by gej

Navigation Processing Finished
Starting Fenix Processing

comment:6 Changed 4 years ago by gej

Fenix Processing

Flight line 1 has been removed as it isn't in the logsheet and before plane aligned in.

had to specify nav2_args_extra in config file for flight lines 2 and 6 as specim_qsub.py couldn't find the files.

found sct values

Flightline FENIX
2 0.96
3 1.10
4 1.01
5 0.98
6 0.97

Nav sync problems appear to be fixed, no other problems to this point

Last edited 4 years ago by gej (previous) (diff)

comment:7 Changed 4 years ago by dac

RCD Processing

Processed RCD files from raw to tif to test calibration file. Used file for 60 mm lens (S/N 27).

comment:8 Changed 4 years ago by stgo

Lidar processing

Starting processing

comment:9 Changed 4 years ago by stgo

Lidar Processing

There's no overlap in the Lidar data and the files are very small. Processing full waveform.

comment:10 Changed 4 years ago by stgo

Lidar Processing

Ready for delivery check

comment:11 Changed 4 years ago by gej

Lidar Delivery check

No problems found.

Ready for delivery

comment:12 Changed 4 years ago by asm

Lidar Delivery

Delivered via FTP (arsf19). Notification has been sent to PI.

comment:13 Changed 4 years ago by gej

Hyperspectral Processing
Fenix processed using new calibration.
Have created delivery and readme.

Ready for Delivery check

comment:14 Changed 4 years ago by asm

Hyperspectral Delivery Check

Found some problems with flightline 03, gej to check again SCT values.

comment:15 Changed 4 years ago by gej

Hyperspectral processing

Found correct SCT for flight 3
SCT is now 1.00

Creating mapped file now.

comment:16 Changed 4 years ago by gej

Hyperspectral Processing

flightline 3 has been remapped, screenshots, mosaic and readme created.

Ready for Re-DC

comment:17 Changed 4 years ago by asm

Fenix DC

-All flightlines seems to have correct SCT values now.
-Checked green vegetation spectra with Py6S. Good match for all flightlines.
-Tested all tif files created from apl commands.
-Corrected new line starting with lowercase on readme and changed slightly the sentence.
-All checks made, no further problems ecountered.

A check on spectra will be done by lah, after that will zip files and project will be ready to go.

comment:18 Changed 4 years ago by lah

Fenix DC - spectra
Spectra for all lines look good. There is a narrow spike at the end of the VNIR region which is quite intense for the start of line 2 where there is a lower signal level, but this is well within the mask region. Spectral offset looks to be well within the 5nm spectral resolution of the SWIR bands, and generally looks better than Malaysian data.

Would be nice to keep the Py6S outputs for spectra checking. Don't forget to delete tmp_cmd_check though.

Ready to deliver once files have finished zipping.

comment:19 Changed 4 years ago by asm

Fenix delivery

Zipped files has been created. Fenix delivery is ready to go.

comment:20 Changed 4 years ago by asm

Fenix Delivery

Delivered via FTP (arsf19). Notification has been sent to PI.

comment:21 Changed 4 years ago by gej

RCD processing

Tiff images tagged. Creating delivery

comment:22 Changed 4 years ago by gej

RCD processing

Delivery and Readme created, ready for delivery check

comment:23 Changed 4 years ago by dac

RCD Delivery Check

Starting delivery check.

comment:24 Changed 4 years ago by dac

RCD Delivery Check

All looks fine - ready to deliver.

comment:25 Changed 4 years ago by dac

RCD Delivery

Sent to PI via FTP (arsf19) 17/12/2015

comment:26 Changed 4 years ago by stgo

LiDAR processing

An issue was discovered with the scan angle filtering, the lidar data has been reprocessed with a wider scan angle of 45 degrees. The wider scanline size has created an overlap region between lines 2 and 3 and lines 3 and 4, this has not introduced a pitch or roll error.

I have created a new delivery which is ready for check.

comment:27 Changed 4 years ago by gej

Starting Lidar Delivery Check

comment:28 Changed 4 years ago by gej

Lidar deliver check

All looks fine,

Ready to deliver

comment:29 Changed 4 years ago by dac

LiDAR data delivery

Redelivered via FTP (arsf19) - 26/02/2016

comment:30 Changed 3 years ago by mark1

Fenix zipped mapped line 03 is corrupt and won't unzip - this file should be re-created before archiving.

comment:31 Changed 3 years ago by dac

Looks like this line was modified after being delivered. Not sure how this happened. Have started reprocessing.

comment:32 Changed 3 years ago by dac

Reprocessed line 3 and replaced mapped zipped file.

comment:33 Changed 3 years ago by dac

Archiving

Started uploading to NEODC.

comment:34 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Data now available from NEODC. Closing ticket.

Note: See TracTickets for help on using tickets.